OpenStudy (anonymous):

Please solve!!!! best answer gets medal. Solve. 2x2 + 3x = -8

OpenStudy (tkhunny):

Have you considered the Quadratic Formula or Completing the Square?

OpenStudy (anonymous):

quadratic formula

OpenStudy (pixiedust1):

Do you want explanation or just answer?

OpenStudy (anonymous):

i know the answer but i need to know how to solve it

OpenStudy (pixiedust1):

2*2+3*-4=-8

OpenStudy (pixiedust1):

Punch that into a calculator

OpenStudy (pixiedust1):

BAMB!! Solved :D

OpenStudy (anonymous):

The second one is the answer the quantity of negative 3 plus or minus 11 i all over 4 the quantity of negative 3 plus or minus i square root 55 all over 4 negative 3 plus or minus i square root 23 3 plus or minus i square root 11

OpenStudy (pixiedust1):

Huh?

OpenStudy (anonymous):

that isnt a answer choice, i got -5

OpenStudy (anonymous):

Is your problem is\[2x^2+3x=-8\]The next step is to move everything to one side.\[2x^2+3x+8=0\] Now we can utilize the Quadratic formula: \[-b \pm \frac{ \sqrt{b^2-4ac} }{ 2a }\] plug our values \[-3 \pm \frac{ \sqrt{3^2-4(2)(8)} }{ 2(2) }\] Once we solve we get \[-3 \pm \frac{ \sqrt{-55} }{ 4 }\] Since we can't take the square root of a negative number, recall that\[i=\sqrt{-1}\] so we can can take it out of our negative root to get. \[-3 \pm \frac{ i\sqrt{55} }{ 4 }\]